# SatFusion: A Unified Framework for Enhancing Satellite IoT Images via Multi-Temporal and Multi-Source Data Fusion

### v1

@@ 2.1. Multi-Temporal Data Fusion
$2.$ Multi-Source Data Fusion
@@
Despite these advances, existing pansharpening networks still face key limitations in real-world Sat-IoT settings. First, most methods rely on a single LRMS and single HRPAN input pair, lacking the ability to exploit abundant multi-temporal observations. Second, since the pioneering PNN model (Masi et al., 2016), nearly all subsequent works have followed the Wald protocol for dataset synthesis, which assumes perfect spatial alignment and noise-free conditions (Wald et al., 1997). This setting diverges from practical Sat-IoT scenarios, where misalignment, blur, and sensor noise are common. Finally, as shown in Figure  2(b), most pansharpening networks first interpolate the input lrmslrms image to match the spatial dimensions of the PanPan image before passing them to the fusion module. The fused features are then processed through either residual connections (Figure  2(d)) or direct outputs (Figure  2(c)) to produce SRSR. Such dependence on a single interpolated input makes these models highly sensitive to LRMS image quality, further limiting robustness under real-world acquisition variability.
@@ 4.2. Implementation Details of Core Modules
where the output super-resolved feature $SR_{MTIF}\in\mathbb{R}^{H\times W\times C_{ms}}$}. The function of the decoder is to adjust the spatial feature dimensions of the image so that it aligns with the input HRPAN image PanPan; and to adjust the channel feature dimension of the image so that it matches the target image. We implement the decoder using a sub-pixel convolution (PixelShuffle) block:
@@
where $x^{PixelShuffleBlock}\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times C_{ms}}$ is the feature map output by the PixelShuffle Block, and $Conv2d$ adjusts the channel dimension to be divisible by ${\gamma}^{2}$, ensuring the correct operation of the sub-pixel convolution module. The sub-pixel convolution increases the spatial resolution of the feature maps by rearranging pixels from the channel dimension into the spatial dimension. When data enter the PixelShuffle Block, pixel rearrangement is performed first, followed by a stack of Conv, Batch Normalization, and PReLU functions to obtain the output. $x^{PixelShuffleBlock}$ is optionally processed by the ResizeResize module:
@@
where ResizeResize adjusts the spatial dimensions by interpolation to ensure that the output features of the MTIF module are spatially aligned with the input PanPan.
@@
The MSIF module is composed of a pansharpening module. By feeding $SR_{MTIF}$ and PanPan into the pansharpening module, the fused features $SR_{MSIF}$ are obtained:
@@
Here, the fused features $SR_{MSIF}\in\mathbb{R}^{H\times W\times C_{ms}}$, and $f_{Sharpen}$ represents the fusion module of the pansharpening network designed by previous researchers (Sharpen module, as shown in Figure 2(c)(d)). The key difference from prior work lies in the fact that we take the multi-temporal fused features $SR_{MSIF}$ together with PanPan as inputs to the sharpen module, where the spatial dimensions of $SR_{MSIF}$ and PanPan are already aligned, rather than merely upsampling a single lrmslrms band by interpolation to match the resolution of PanPan.
Fusion Composition Module. Its goal is to fuse the outputs of the MTIF module and the MSIF module, so as to adaptively exploit the advantages of both multi-temporal and multi-source image fusion. The inputs are the output $SR_{MTIF}$ from the MTIF module and the output $SR_{MSIF}$ from the MSIF module, and the output is the final single HRMS image SRSR. The Fusion composition module consists of residual connections and convolution adjustment operations, and finally outputs SRSR:
@@
Here, the output is the HRMS image $SR\in\mathbb{R}^{H\times W\times C_{ms}}$, and ConvConv denotes a 2D convolution with kernel size 111*1. The 111*1 convolution plays a role in adjusting spectral characteristics (Razzak et al., 2023). By placing the convolution operation after the residual connection, the model is encouraged to learn spectral features more effectively.
@@ 4.3. Loss Function
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are the corresponding weights, satisfying $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=1$. SRSR denotes the image output by SatFusion, and GTGT denotes the ground truth image. $\mathcal{L}_{\mathrm{MAE}}$ is the $l_{1}$ loss, and $\mathcal{L}_{\mathrm{MSE}}$ is the $l_{2}$ loss, both of which are commonly used to supervise the outputs of MISR and pansharpening networks. Prior works retained the same loss functions while modifying model architectures to ensure fair comparison. On top of these loss functions, we introduce new supervisory signals: $\mathcal{L}_{\mathrm{SSIM}}$ is a texture loss function focusing on the learning of spatial features, while $\mathcal{L}_{\mathrm{SAM}}$ is a spectral loss function emphasizing the learning of spectral features.
Since discrepancies exist between the acquisition conditions of input images and the ground truth, SRSR and GTGT may exhibit horizontal/vertical shifts as well as overall brightness deviations. Following prior work, we perform brightness adjustment and boundary cropping before computing the loss in some experiments (An et al., 2022; Deudon et al., 2020; Salvetti et al., 2020). For more details on SatFusion’s loss calculation, please refer to Appendix D. It is worth emphasizing that we ensure consistency in loss calculation during experiments. Specifically, when comparing SatFusion with baseline methods, we apply the same adjustment operations to SRSR before computing the loss, and use the same functions for loss computation across the entire training, validation, and testing pipeline. This guarantees that improvements in evaluation metrics directly demonstrate the effectiveness of the architecture.
@@ 5.1. Experimental Setup
We conduct experiments on both the real dataset and synthetic datasets constructed using the Wald protocol. It is worth noting that SatFusion takes multiple multi-temporal LRMS images and a single HRPAN image as input. However, as shown in Figure 4(left), prior pansharpening works use Wald protocol to construct synthetic lrmslrms and PanPan images as input, which has inherent idealized limitations (Section 2.2). As shown in Figure 4(right), we introduce pixel shifts and noise perturbations into the Wald protocol to construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$, which are then combined with the PanPan image as the input of SatFusion. Experiments are conducted on both real datasets and synthetic datasets with added noise perturbations, making them more consistent with the data conditions in Sat-IoT scenarios.
@@ 5.2. Main Experiments
On the real dataset, we validate that the SatFusion framework is beneficial for enhancing both MISR and Pansharpening networks. We conduct experiments on the original WorldStrat dataset, where SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and PanPan as inputs,   the MISR network takes $\{lrms_{t}\}_{t=1}^{T=8}$ as input, and the Pansharpening network takes a randomly selected single lrmslrms and PanPan as inputs.
@@
On the synthetic dataset, we validate that the SatFusion framework is beneficial for enhancing Pansharpening networks. Using the WorldStrat, WV3, QB, and GF2 datasets, we construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$ with pixel shifts and noise disturbances introduced through the Wald protocol. SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and PanPan as inputs, while the Pansharpening network takes a randomly selected lrmslrms and PanPan as inputs.
As shown in Table 2 and Table 1(b), SatFusion significantly improves the quality of the fused images on the test sets of the synthetic datasets. We visualize the fused images and the error maps between the SRSR and the GTGT in Figure 6, verifying its ability to adaptively fuse multi-source features in scenarios with complex disturbances. This further demonstrates the effectiveness and advantages of SatFusion in integrating MISR and Pansharpening networks.
@@ 6.1. Practical Significance of SatFusion
Assuming each pixel in every channel occupies one unit of storage space, let NN denote the number of locations. Then the input data size is $D_{\text{input}}=N\times\Big(T\cdot\tfrac{H}{\gamma}\cdot\tfrac{W}{\gamma}\cdot C_{ms}\;+\;H\cdot W\cdot 1\Big),$ and the output data size is $D_{\text{output}}=N\times(H\cdot W\cdot C_{ms}).$ While redundancy reduction is guaranteed, the total physical data volume depends on specific conditions. Because the fused output typically has higher spatial resolution and more channels, the enhanced information richness may increase total data size. Only when $T$ is sufficiently large does the compression benefit outweigh this increase, satisfying $D_{\text{output}}<D_{\text{input}}$. As shown in Figure 7, this advantage becomes increasingly pronounced in Sat-IoT scenarios characterized by abundant multi-temporal data.
@@ 6.2. Future Prospects
SatFusion takes $\{lrms_{t}\}_{t=1}^{T}$ and Pan{Pan} as inputs, where the multi-temporal $\{lrms_{t}\}_{t=1}^{T}$ images not only provide complementary information but also contribute to learning deep features aligned with PanPan. However, in real-world scenarios, multi-temporal PAN images also exist but were not utilized in our design. Intuitively, incorporating multi-temporal PAN images through novel architectural designs could further improve fusion performance.
@@ Appendix A Hyperparameter Experiments
We also vary the super-resolution scaling factor $\gamma$ on the original WorldStrat dataset. As shown in Table 3, SatFusion achieves the best performance at scaling factors $\gamma=2,5$ (In Table  1, $\gamma\approx 3$). Moreover, for relatively small scaling factors, the fused super-resolved MS images SRSR produced by SatFusion exhibit more stable quality.
@@ Appendix D Details of Loss Computation
To handle potential image misalignment, SatFusion crops both GTGT and SRSR to compensate for spatial shifts and returns the minimum loss value (Cornebise et al., 2022). Cropping is implemented using a Lanczos convolution kernel. Lanczos is a high-quality image resampling filter based on the sinc function. It achieves a balance between computational efficiency and anti-aliasing by truncating the infinite support of the sinc function with a window function, thereby reducing ringing artifacts while preserving sharpness. The initial offset $x$ is generated from $p_{x}$ (absolute shift), stepstep (shift stride), and $a$ (kernel lobes):
@@
Here, $patch(\cdot)$ is a cropping function that adjusts the image size to match $GT_{i}^{\prime}$. LossLoss refers to the composite weighted loss introduced in Section 4.3, with weights $\lambda_{1}=0.3$, $\lambda_{2}=0.3$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.2$.

### v2

@@ 2.1. Multi-Temporal Data Fusion
2.2. Multi-Source Data Fusion
@@
Despite these advances, existing pansharpening networks still face key limitations in real-world Sat-IoT settings. First, most methods rely on a single LRMS and single HRPAN input pair, lacking the ability to exploit abundant multi-temporal observations. Second, since the pioneering PNN model (Masi et al., 2016), nearly all subsequent works have followed the Wald protocol for dataset synthesis, which assumes perfect spatial alignment and noise-free conditions (Wald et al., 1997). This setting diverges from practical Sat-IoT scenarios, where misalignment, blur, and sensor noise are common. Finally, as shown in Figure  2(b), most pansharpening networks first interpolate the input lrmslrms image to match the spatial dimensions of the $Pan$ image before passing them to the fusion module. The fused features are then processed through either residual connections (Figure  2(d)) or direct outputs (Figure  2(c)) to produce $SR$. Such dependence on a single interpolated input makes these models highly sensitive to LRMS image quality, further limiting robustness under real-world acquisition variability.
@@ 4.2. Implementation Details of Core Modules
where the output super-resolved feature $SR_{MTIF}\in\mathbb{R}^{H\times W\times C_{ms}}$}. The function of the decoder is to adjust the spatial feature dimensions of the image so that it aligns with the input HRPAN image $Pan$; and to adjust the channel feature dimension of the image so that it matches the target image. We implement the decoder using a sub-pixel convolution (PixelShuffle) block:
@@
where $x^{PixelShuffleBlock}\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times C_{ms}}$ is the feature map output by the PixelShuffle Block, and $Conv2d$ adjusts the channel dimension to be divisible by ${\gamma}^{2}$, ensuring the correct operation of the sub-pixel convolution module. The sub-pixel convolution increases the spatial resolution of the feature maps by rearranging pixels from the channel dimension into the spatial dimension. When data enter the PixelShuffle Block, pixel rearrangement is performed first, followed by a stack of Conv, Batch Normalization, and PReLU functions to obtain the output. $x^{PixelShuffleBlock}$ is optionally processed by the $Resize$ module:
@@
where $Resize$ adjusts the spatial dimensions by interpolation to ensure that the output features of the MTIF module are spatially aligned with the input $Pan$.
@@
The MSIF module is composed of a pansharpening module. By feeding $SR_{MTIF}$ and $Pan$ into the pansharpening module, the fused features $SR_{MSIF}$ are obtained:
@@
Here, the fused features $SR_{MSIF}\in\mathbb{R}^{H\times W\times C_{ms}}$, and $f_{Sharpen}$ represents the fusion module of the pansharpening network designed by previous researchers (Sharpen module, as shown in Figure 2(c)(d)). The key difference from prior work lies in the fact that we take the multi-temporal fused features $SR_{MSIF}$ together with $Pan$ as inputs to the sharpen module, where the spatial dimensions of $SR_{MSIF}$ and $Pan$ are already aligned, rather than merely upsampling a single lrmslrms band by interpolation to match the resolution of $Pan$.
Fusion Composition Module. Its goal is to fuse the outputs of the MTIF module and the MSIF module, so as to adaptively exploit the advantages of both multi-temporal and multi-source image fusion. The inputs are the output $SR_{MTIF}$ from the MTIF module and the output $SR_{MSIF}$ from the MSIF module, and the output is the final single HRMS image $SR$. The Fusion composition module consists of residual connections and convolution adjustment operations, and finally outputs $SR$:
@@
Here, the output is the HRMS image $SR\in\mathbb{R}^{H\times W\times C_{ms}}$, and $Conv$ denotes a 2D convolution with kernel size 111*1. The 111*1 convolution plays a role in adjusting spectral characteristics (Razzak et al., 2023). By placing the convolution operation after the residual connection, the model is encouraged to learn spectral features more effectively.
@@ 4.3. Loss Function
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are the corresponding weights, satisfying $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=1$. $SR$ denotes the image output by SatFusion, and $GT$ denotes the ground truth image. $\mathcal{L}_{\mathrm{MAE}}$ is the $l_{1}$ loss, and $\mathcal{L}_{\mathrm{MSE}}$ is the $l_{2}$ loss, both of which are commonly used to supervise the outputs of MISR and pansharpening networks. Prior works retained the same loss functions while modifying model architectures to ensure fair comparison. On top of these loss functions, we introduce new supervisory signals: $\mathcal{L}_{\mathrm{SSIM}}$ is a texture loss function focusing on the learning of spatial features, while $\mathcal{L}_{\mathrm{SAM}}$ is a spectral loss function emphasizing the learning of spectral features.
Since discrepancies exist between the acquisition conditions of input images and the ground truth, $SR$ and $GT$ may exhibit horizontal/vertical shifts as well as overall brightness deviations. Following prior work, we perform brightness adjustment and boundary cropping before computing the loss in some experiments (An et al., 2022; Deudon et al., 2020; Salvetti et al., 2020). For more details on SatFusion’s loss calculation, please refer to Appendix D. It is worth emphasizing that we ensure consistency in loss calculation during experiments. Specifically, when comparing SatFusion with baseline methods, we apply the same adjustment operations to $SR$ before computing the loss, and use the same functions for loss computation across the entire training, validation, and testing pipeline. This guarantees that improvements in evaluation metrics directly demonstrate the effectiveness of the architecture.
@@ 5.1. Experimental Setup
We conduct experiments on both the real dataset and synthetic datasets constructed using the Wald protocol. It is worth noting that SatFusion takes multiple multi-temporal LRMS images and a single HRPAN image as input. However, as shown in Figure 4(left), prior pansharpening works use Wald protocol to construct synthetic lrmslrms and $Pan$ images as input, which has inherent idealized limitations (Section 2.2). As shown in Figure 4(right), we introduce pixel shifts and noise perturbations into the Wald protocol to construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$, which are then combined with the $Pan$ image as the input of SatFusion. Experiments are conducted on both real datasets and synthetic datasets with added noise perturbations, making them more consistent with the data conditions in Sat-IoT scenarios.
@@ 5.2. Main Experiments
On the real dataset, we validate that the SatFusion framework is beneficial for enhancing both MISR and Pansharpening networks. We conduct experiments on the original WorldStrat dataset, where SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and $Pan$ as inputs,   the MISR network takes $\{lrms_{t}\}_{t=1}^{T=8}$ as input, and the Pansharpening network takes a randomly selected single lrmslrms and $Pan$ as inputs.
@@
On the synthetic dataset, we validate that the SatFusion framework is beneficial for enhancing Pansharpening networks. Using the WorldStrat, WV3, QB, and GF2 datasets, we construct synthetic multi-temporal images $\{lrms_{t}\}_{t=1}^{T}$ with pixel shifts and noise disturbances introduced through the Wald protocol. SatFusion takes $\{lrms_{t}\}_{t=1}^{T=8}$ and $Pan$ as inputs, while the Pansharpening network takes a randomly selected lrmslrms and $Pan$ as inputs.
As shown in Table 2 and Table 1(b), SatFusion significantly improves the quality of the fused images on the test sets of the synthetic datasets. We visualize the fused images and the error maps between the $SR$ and the $GT$ in Figure 6, verifying its ability to adaptively fuse multi-source features in scenarios with complex disturbances. This further demonstrates the effectiveness and advantages of SatFusion in integrating MISR and Pansharpening networks.
@@ 6.1. Practical Significance of SatFusion
Assuming each pixel in every channel occupies one unit of storage space, let $N$ denote the number of locations. Then the input data size is $D_{\text{input}}=N\times\Big(T\cdot\tfrac{H}{\gamma}\cdot\tfrac{W}{\gamma}\cdot C_{ms}\;+\;H\cdot W\cdot 1\Big),$ and the output data size is $D_{\text{output}}=N\times(H\cdot W\cdot C_{ms}).$ While redundancy reduction is guaranteed, the total physical data volume depends on specific conditions. Because the fused output typically has higher spatial resolution and more channels, the enhanced information richness may increase total data size. Only when $T$ is sufficiently large does the compression benefit outweigh this increase, satisfying $D_{\text{output}}<D_{\text{input}}$. As shown in Figure 7, this advantage becomes increasingly pronounced in Sat-IoT scenarios characterized by abundant multi-temporal data.
@@ 6.2. Future Prospects
SatFusion takes $\{lrms_{t}\}_{t=1}^{T}$ and Pan{Pan} as inputs, where the multi-temporal $\{lrms_{t}\}_{t=1}^{T}$ images not only provide complementary information but also contribute to learning deep features aligned with $Pan$. However, in real-world scenarios, multi-temporal PAN images also exist but were not utilized in our design. Intuitively, incorporating multi-temporal PAN images through novel architectural designs could further improve fusion performance.
@@ Appendix A Hyperparameter Experiments
We also vary the super-resolution scaling factor $\gamma$ on the original WorldStrat dataset. As shown in Table 3, SatFusion achieves the best performance at scaling factors $\gamma=2,5$ (In Table  1, $\gamma\approx 3$). Moreover, for relatively small scaling factors, the fused super-resolved MS images $SR$ produced by SatFusion exhibit more stable quality.
@@ Appendix D Details of Loss Computation
To handle potential image misalignment, SatFusion crops both $GT$ and $SR$ to compensate for spatial shifts and returns the minimum loss value (Cornebise et al., 2022). Cropping is implemented using a Lanczos convolution kernel. Lanczos is a high-quality image resampling filter based on the sinc function. It achieves a balance between computational efficiency and anti-aliasing by truncating the infinite support of the sinc function with a window function, thereby reducing ringing artifacts while preserving sharpness. The initial offset $x$ is generated from $p_{x}$ (absolute shift), stepstep (shift stride), and $a$ (kernel lobes):
@@
Here, $patch(\cdot)$ is a cropping function that adjusts the image size to match $GT_{i}^{\prime}$. $Loss$ refers to the composite weighted loss introduced in Section 4.3, with weights $\lambda_{1}=0.3$, $\lambda_{2}=0.3$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.2$.